OpenStudy (anonymous):

For the right triangle shown, the lengths of two sides are given. Find the third side. Leave your answers in simplified, radical form. Type your answer in the answer box provided. a = 4, b = 4, c =

OpenStudy (anonymous):



OpenStudy (anonymous):

simple Pythagorean theorem: a^2+b^2=c^2 4^2+4^2=c^2 c=sqrt(32) c = \[4\sqrt(2)\]